Your favorite Kalesa moves, with an v0 of 10 m/s, accelerates at a rate of 50 m/s2 for 3s, 


1
Expert's answer
2021-04-19T17:10:18-0400

...then the final velocity is


"v_f=v_i+at=10+50·3=160\text{ m/s}"